OpenStudy (anonymous):

Calculate the derivative of the function. h(x) = 9 / (x^2+x+1)^2

OpenStudy (anonymous):

did you try the quotient rule?

OpenStudy (anonymous):

it's pretty straightforward here...

zepdrix (zepdrix):

Hmm this looks just like the last one chris :) nice problem to apply the power rule on (if we first deal with getting that term to the numerator).

OpenStudy (anonymous):

\[\frac{ f }{ g }'=\frac{ f'g-g'f }{ g^2 }\]

OpenStudy (anonymous):

\[\frac{ -9(x^2+x+1)^2 }{ (x^2+x+1)^4 }\]

OpenStudy (anonymous):

yea i have to use the chain rule again

OpenStudy (anonymous):

i just have a hard time using the chain rule for this one

OpenStudy (anonymous):

no i used qoutient rule

OpenStudy (anonymous):

chain rule will look like\[h(x)=9(x^2+x+1)^{-2}\] \[\frac{ (-2)9(2x+1) }{ (x^2+x+1)^3}\]

OpenStudy (anonymous):

\[\frac{ -9[(x^2+x+1)^2]' }{ (x^2+x+1)^4 }=\frac{ -2(9)(x^2+x+1)(2x+1) }{ (x^2+x+1)^2 }\]finally sorted

OpenStudy (anonymous):

you wouldnt multiply -2(9)?

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

im sorry but i plugged that into my assignment and it was incorrect

OpenStudy (anonymous):

try this\[\frac{ -18(2x+1) }{ (x^2+x+1)^3 }\]

OpenStudy (anonymous):

sorry its still incorrect, are you presenting the right material?

OpenStudy (anonymous):

i wouldn't use the quotient rule for this one because the numerator is a constant just butting in ....

zepdrix (zepdrix):

 Hmmm the answer Jonask gave you looks correct. I'm not sure why it's not accepting your answer. Maybe you're having trouble typing the format in correctly? :o